# Learning to Assist: Physics-Grounded Human-Human Control via Multi-Agent Reinforcement Learning

### 11 Mar 2026 v1

@@ 1 Introduction
We evaluate AssistMimic on the Inter-X [26] and HHI-Assist [19] datasets, both of which feature tightly coupled, support-oriented human–human interactions. Across these benchmarks, AssistMimic is the only method to track closely interacting, force exchanging human motions, achieving substantially higher task success rates (Inter-X: 83%, HHI-Assist: 66%) and demonstrating strong robustness to unseen disturbances such as stumbles and load shifts. Empirically, AssistMimic accurately reproduces reference motions across diverse assistive scenarios—including recipients lying on the floor, seated in chairs, or resting in beds—where the assisting agent learns to provide appropriate physical support and lifting assistance. To objectively assess assistance quality, we reduce the recipient’s PD gains and maximum torque limits to constrain self-stabilization, thereby isolating the contribution of the learned assistive controller.
@@ 3 Method
The overall architecture of AssistMimic is illustrated in Figure  3. It formulates human-human interactive motion imitation as a Multi-Agent Reinforcement Learning (MARL) problem. Our framework focuses on assistive scenarios involving two distinct agents: an assistant who provides physical support, and a recipient who requires assistance. To simulate physical impairment, we impose specific constraints on the recipient’s dynamic parameters, including reduced maximum joint torques and lowered Proportional-Derivative (PD) gains ($k_{p},k_{d}$). Both agents learn tracking policies that are optimized for reference motion fidelity under these physics constraints. Each policy perceives the environment through a partner-aware, ego-centric state space. To bootstrap learning, we initialize both agents using a pre-trained motion prior, allowing them to inherit common coordination skills before trained to acquire role-specific behaviors.
@@ 3.1 Problem Formulation
For the recipient ($m=\mathrm{R}$), the task reward consists of tracking, power, and assist stability terms; the full decomposition is provided in D.1. However, for the supporter ($m=\mathrm{S}$), we modify this objective to satisfy interaction requirements. First, as detailed in Section 3.3, the supporter’s hand reference trajectories are dynamically retargeted relative to the recipient’s body to maintain valid relative positioning. Second, as described in Section 3.4, the standard hand-tracking reward terms are replaced by contact-incentivizing objectives when in close proximity to the recipient, prioritizing functional support over strict kinematic adherence.
@@ 3.4 Contact-Promoting Rewards
where $\alpha,\beta$ are hyperparameter for scaling, $b_{\mathrm{contact}}$ is a sparse bonus awarded upon establishing contact, and $f_{i,t}$ is the measured finger-contact forces aggregated with a safety-aware soft saturation function:
@@
We cluster motions based on subject ids, and train a single “specialist” policy per cluster with PPO [21]. We distill specialist policies into a “generalist” policy to be able to handle more diverse reference motions, using DAgger [18]. For both specialist and generalist training, we apply tight early termination [24] with a 0.25 m pose-deviation threshold. For initialization, we adopt Physical State Initialization (PSI [30]): instead of sampling from noisy reference trajectories that often contain penetrations, we seed episodes from states collected in recent rollouts.
@@ 4 Experiments
  • Inter-X [26] This is the largest two-person HHI dataset and stores motions in the SMPL-X [14] format. Unlike prior work that focuses only on contact-less social interactions, such as high-five [9, 6], we extracted 60 motions (30 paired motions) from sequences labeled “Help-up” to cover a wide range of assistive strategies. These motions exhibit high diversity in both spatial approach and contact topology. They cover various approach directions (e.g., front, back) and distinct support modalities, ranging from single-arm grasping to dual-hand shoulder support.
  • HHI-Assist [19] This dataset captures caregiving motions on a bed or chair. It includes specialized actions such as supporting the recipient’s shoulder while lifting them from the bed, or stabilizing the knees and rotating the body to a seated posture. We use 10 representative caregiving motion clips (details in Appendix C).
Evaluation settings We evaluate AssistMimic and its baselines under four scenarios:
  1. (i)
    Specialist policy evaluation. Both our method and the baselines are trained on approximately 8–10 motion clips, grouped by subject, to assess performance in a specialist setting.
  2. (ii)
    Generalist policy evaluation. As Inter-X includes a broad range of get-up and support strategies, we train and evaluate a single policy on 30 diverse clips to test whether AssistMimic can reproduce a wide variety of support behaviors with one unified controller.
  3. (iii)
    Support under unseen dynamics. We assess robustness to recipient dynamics not seen during training by modifying physical parameters at test time: halving the PD gains from the training phase, increasing body mass to $1.2\times$ or $1.5\times$, and reducing maximum hip torque to $0.5\times$ the training value.
  4. (iv)
    Tracking generated kinematic trajectories. We apply our tracking policy to outputs from a generative model that synthesizes two-person interactions and evaluate the quality of tracking on these unseen, generated motion sequences.
@@
Table 4: Recipient COM standard deviation ($\downarrow$)
Model Seen Mass $\times 1.5$ Max hip $\tau$ $\times 0.5$
Ours 0.0921 0.0738 0.0865
(-) Dyn Retar 0.1038 0.0902 0.0924
(-) Cont 0.0938 0.0838 0.0849
@@
  • Mean Per Joint Position Error (MPJPE), which measures the average positional deviation between the simulated and reference joint positions for both the supporter and the recipient.
  • COM Stability, specifically for the HHI-Assist dataset where steady support on a bed is required. We calculate the standard deviation of the recipient’s Center of Mass (COM) position throughout the episode. A lower standard deviation reflects steadier, more controlled assistance. To ensure a consistent number of frames, this metric is computed exclusively for motion sequence subsets where all compared models achieved success.
@@ 4.1 Specialist Policy Evaluation
We first qualitatively demonstrate that the recipient kinematic replay paradigm is ill-posed for the assistive tasks. As illustrated in Figure 4 (left), applying the kinematic replay approach used in Human-X [6] to our scenario results in the recipient “standing up” independently according to a pre-defined trajectory, regardless of the presence of support (indicated by yellow arrows). Consequently, the assistant’s actions exert no physical influence on the partner’s state, preventing the model from learning functional support behaviors. Furthermore, in Figure 4 (right), the kinematically fixed poses cause severe interpenetration; the physics engine then triggers massive restorative forces to resolve these overlaps, which eventually blow the characters apart. These failures confirm that a kinematic-replay paradigm is fundamentally inapplicable to the recipient in our task.
Following these qualitative observations, we provide a rigorous quantitative evaluation to demonstrate how our joint multi-agent formulation overcomes these limitations. Tables 2 and 3 compare the specialist policies of AssistMimic against its ablative variants. Table 2 also shows the accuracy of Frozen-Recipient (decoupled sequential learning) approach. Qualitative results on the HHI-Assist dataset are also presented in Figure 5. Our analysis yields the following key insights:
  • Joint Training vs. Sequential Learning: As shown in Table 2, the Sequential Training baseline is substantially outperformed by our joint MARL formulation on the Inter-X dataset. These results align with the premise of our work: in high-contact assistive tasks, although the recipient is incapable of performing the task independently, it is crucial for the recipient to actively learn how to coordinate and receive support, rather than merely replaying a motion. This bidirectional adaptation is the key factor enabling AssistMimic to handle such complex, force-exchanging interactions within physically natural motions.
  • Importance of Motion Prior Initialization: Bootstrapping from the single-agent motion prior is indispensable for convergence. Without this initialization, the RL process fails completely (0% success rate) or succumbs to reward hacking—exploiting the objective function to produce unnatural, non-functional motions. (Note: Entries marked with in Table 3 denote these degenerate reward-hacking failures).
@@
  • Enhanced Stability in Bed-based Assistance: As summarized in Table 4, AssistMimic achieves lower recipient COM standard deviation compared to the ablated variants on the HHI-Assist when evaluated on motions successfully completed by all policies. Together with the results in Table 3, this indicates that our approach achieves both high success rates and stable assistance.
Figure 5: Qualitative evaluation with HHI-Assist. The red boxes highlight whether the supporter’s hands correctly adjust to the recipient’s position and provide appropriate support.
@@ 4.2 Generalist Policy evaluation
Table 5 summarizes the tracking accuracy for generalist policies evaluated on the Inter-X dataset. When trained directly on a diverse subset of 30 interaction clips, AssistMimic achieves a success rate of 39.8%. This represents a breakthrough for physics-based control in this domain; to the best of our knowledge, AssistMimic is the first method to succeed in such a diverse set of tightly coupled interactive motions. Furthermore, we observe that distilling specialist policies into a single generalist agent via DAgger yields substantial performance gains. This distillation process boosts the success rate to 64.7%, establishing a new benchmark for diverse, close-contact interactive motion synthesis. The top two rows of Figure LABEL:fig:teaser present qualitative results on the Inter-X dataset obtained with a single generalist policy trained with AssistMimic. For additional qualitative results, please see Appendix B.
@@ 4.4 Tracking of Generated Interactions
Figure 6 illustrates the results of applying AssistMimic to trajectories synthesized by a text-conditioned motion diffusion model (details in Appendix E). The generative model employs an MDM-style [25] denoising transformer architecture to auto-regressively generate joint positions for both the supporter and recipient, conditioned on text descriptions. The diffusion model was trained on ”Help-up” interactions from the Inter-X dataset. During inference, we prompt it with unseen text descriptions from the same category. While the resulting kinematic outputs capture the semantic structure of the interaction, they frequently contain physical inconsistencies typical of diffusion-generated motion (e.g., foot sliding, penetration). AssistMimic successfully converts dense interaction kinematics into physically plausible motions with high accuracy. This result demonstrates that the AssistMimic framework is effective not only for real interaction data but also for generated data, highlighting its generality and broad applicability.
@@ 4.5 Generalization to Unseen Supportive Actions
Figure 7: Tracking results on unseen actions.
As shown in Figure 7, AssistMimic also generalizes to unseen interaction categories from Inter-X, such as support-by-arm. Even though this interaction pattern is not observed during training, the humanoid successfully follows the motion while placing a supportive hand on the partner’s arm and avoiding body collisions.
Failure cases Figure 8 illustrates typical failure modes, particularly in motions where the supporter must grasp and lift the recipient’s arms. These failures stem primarily from limited hand dexterity: accurately grasping a human arm requires precise, coordinated finger movements, which are difficult to learn from noisy demonstrations. The substantial weight of the recipient’s body (approximately 50 kg) further compounds the challenge, making stable lifting difficult.
4.6 Limitations
AssistMimic has several limitations that point toward promising directions for future work. First, a tighter integration between the motion planner and the tracking controller could enable more adaptive, real-time coordination between high-level planning and low-level control. second, incorporating visual observations of the partners’ states and transferring the learned behaviors from simulation to real humanoid systems remain key challenges and clear avenues for future research.
Figure 8: Failure cases.
@@ 5 Conclusion
We introduced AssistMimic, a multi-agent reinforcement learning framework for learning physics-aware, tracking-based controllers that reproduce—and adapt to—human–human interactive motions, including assistive behaviors. In contrast to prior approaches that treat interaction partners independently or rely on fixed replayed trajectories, AssistMimic jointly learns policies for both agents, enabling reactive and physically consistent coordination. AssistMimic improves exploration and stability in RL through three key innovations: (1) weight initialization from single-human tracking controllers, which stabilizes early learning; (2) dynamic reference retargeting, allowing the policy to adjust to the recipient’s current pose; and (3) contact-promoting rewards, encouraging physically meaningful interactions. Experiments on the Inter-X and HHI-Assist benchmarks show that AssistMimic is the first method capable of successfully tracking assistive interaction sequences. More broadly, our results demonstrate a generalizable framework for extending single-agent imitation learning to physically grounded multi-agent domains, with promising implications for assistive robotics, collaborative manipulation, and embodied simulation of social interactions.
@@ Appendix B Additional Qualitative comparison
Figure 9 visualizes the ablation results of the specialist policies on the Inter-X dataset. Each row corresponds to a different variant, and frames are shown from left to right in chronological order. The first and second rows compare the best AssistMimic policy (w/o dynamic reference retergeting) and the variant without the contact-promoting reward on the same help-up motion. The third row shows two separate examples for the variant without weight initialization, because the humanoids quickly lose balance and the episodes terminate early.
@@
Appendix C Additional implementation details
@@
For the multi-agent RL optimization, we use Proximal Policy Optimization (PPO) [21] for both the supporter and the recipient policies. Each specialist policy is trained for 2k PPO iterations. The learning rate for both agents is set to $5\times 10^{-6}$, and we apply a step decay by a factor of $0.1$ once training reaches $600$ PPO iterations. The critic network is implemented as an MLP that takes the same input as the actor, augmented with a role label indicating whether the agent is the recipient or the caregiver (supporter).
@@
In many HHI-Assist bed sequences, the supporter steps away from the recipient near the end of the motion after the assistance is completed. Because the recipient is no longer supported during this phase, falls occurring in this stage should not be considered assistive failures. Thus, for quantitative evaluations (Figure 3, 4), we discard the final 15 frames of each 30 FPS motion clip and compute all metrics on the remaining frames.
@@
Appendix D Detailed reward breakdown

### 10 Apr 2026 v2

@@ 1 Introduction
We evaluate AssistMimic on the Inter-X [26] and HHI-Assist [19] datasets, both of which feature tightly coupled, support-oriented human–human interactions. Across these benchmarks, AssistMimic is the only method to track closely interacting, force exchanging human motions, achieving substantially higher task success rates (Inter-X: 83%, HHI-Assist: 73%) and demonstrating strong robustness to unseen disturbances such as stumbles and load shifts. Empirically, AssistMimic accurately reproduces reference motions across diverse assistive scenarios—including recipients lying on the floor, seated in chairs, or resting in beds—where the assisting agent learns to provide appropriate physical support and lifting assistance. To objectively assess assistance quality, we reduce the recipient’s PD gains and maximum torque limits to constrain self-stabilization, thereby isolating the contribution of the learned assistive controller.
@@ 3 Method
The overall architecture of AssistMimic is illustrated in Figure  3. It formulates human-human interactive motion imitation as a Multi-Agent Reinforcement Learning (MARL) problem. Our framework focuses on assistive scenarios involving two distinct agents: an assistant who provides physical support, and a recipient who requires assistance. To simulate physical impairment, we impose specific constraints on the recipient’s dynamic parameters, including reduced maximum joint torques and lowered Proportional-Derivative (PD) gains ($k_{p},k_{d}$). Both agents learn tracking policies that are optimized for reference motion fidelity under these physics constraints. Each policy perceives the environment through a partner-aware, ego-centric state space. To bootstrap learning, we initialize both agents using a pre-trained motion prior, allowing them to inherit common coordination skills before being trained to acquire role-specific behaviors.
@@ 3.1 Problem Formulation
For the recipient ($m=\mathrm{R}$), the task reward consists of tracking, power, and assist stability terms; the full decomposition is provided in Appendix E. However, for the supporter ($m=\mathrm{S}$), we modify this objective to satisfy interaction requirements. First, as detailed in Section 3.3, the supporter’s hand reference trajectories are dynamically retargeted relative to the recipient’s body to maintain valid relative positioning. Second, as described in Section 3.4, the standard hand-tracking reward terms are replaced by contact-incentivizing objectives when in close proximity to the recipient, prioritizing functional support over strict kinematic adherence.
@@ 3.4 Contact-Promoting Rewards
where $\alpha,\beta$ are hyperparameters for scaling, $b_{\mathrm{contact}}$ is a sparse bonus awarded upon establishing contact, and $f_{i,t}$ is the measured finger-contact forces aggregated with a safety-aware soft saturation function:
@@
We cluster motions based on subject ids, and train a single “specialist” policy per cluster with PPO [21]. We distill specialist policies into a “generalist” policy to be able to handle more diverse reference motions, using DAgger [18]. For both specialist and generalist training, we apply tight early termination [24] with a 0.25 m pose-deviation threshold. For initialization, we adopt Physical State Initialization (PSI [30]): instead of sampling from noisy reference trajectories that often contain penetrations, we seed episodes from states collected in recent rollouts. Other hyperparameters are detailed in Appendix D.
@@ 4 Experiments
  • Inter-X [26] This is the largest two-person HHI dataset and stores motions in the SMPL-X [14] format. Unlike prior work that focuses only on contact-less social interactions, such as high-five [9, 6], we extracted 30 paired “Help-up” motions to cover a wide range of assistive strategies. These motions are diverse in both approach and contact patterns, spanning directions such as front and back, and support types from single-arm grasping to two-hand shoulder support.
  • HHI-Assist [19] This dataset captures caregiving motions on a bed or chair. It includes specialized actions such as supporting the recipient’s shoulder while lifting them from the bed, or stabilizing the knees and rotating the body to a seated posture. We use 10 representative caregiving motion clips (details in Appendix D).
Evaluation settings We evaluate AssistMimic and its baselines under four scenarios: (i) Specialist policy evaluation. Both our method and the baselines are trained on approximately 8–10 motion clips, grouped by subject, to assess performance in a specialist setting. (ii) Generalist policy evaluation. As Inter-X includes a broad range of get-up and support strategies, we train and evaluate a single policy on 30 diverse clips to test whether AssistMimic can reproduce a wide variety of support behaviors with one unified controller. (iii) Support under unseen dynamics. We assess robustness to recipient dynamics not seen during training by modifying physical parameters at test time: halving the PD gains from the training phase, increasing body mass to $1.2\times$ or $1.5\times$, and reducing maximum hip torque to $0.5\times$ the training value. (iv) Tracking generated kinematic trajectories. We apply our tracking policy to outputs from a generative model that synthesizes two-person interactions.
@@
  • Mean Per Joint Position Error (MPJPE), which measures the average positional deviation between the simulated and reference joint positions for both agents.
  • COM Stability on HHI-Assist, which requires steady bed support. We measure the temporal stability of the recipient’s center of mass (COM). See Appendix C for details.
@@ 4.1 Specialist Policy Evaluation
We first show that the recipient kinematic replay is ill-posed for assistive tasks. In Figure 4 (left), applying the kinematic replay used in Human-X [6] to our scenario causes the recipient to “stand up” independently along a pre-defined trajectory, regardless of the presence of support (yellow arrows). Consequently, the assistant’s actions do not meaningfully affect the partner’s state, preventing the emergence of functional support behaviors. In Figure 4 (right), the kinematically fixed poses also cause severe interpenetration, which induces large restorative forces in the physics engine and eventually pushes the characters apart. These failures show that kinematic replay is unsuitable for the recipient in our task.
We next demonstrate how our joint MARL formulation addresses these limitations. Tables 2 and 3 compare the specialist policies of AssistMimic against its ablative variants. Table 2 also includes the Frozen-Recipient (decoupled sequential learning) approach. Qualitative results on the HHI-Assist dataset are shown in Figure 5. Our analysis yields the following key insights:
  • Joint Training vs. Sequential Learning: As shown in Table 2, the Sequential Training baseline is substantially outperformed by our joint MARL formulation on the Inter-X dataset. These results align with the premise of our work: in high-contact assistive tasks, although the recipient is incapable of performing the task independently, it is crucial for the recipient to actively learn how to coordinate and receive support, rather than merely replaying a motion. This bidirectional adaptation enables AssistMimic to handle such complex, force-exchanging interactions within physically natural motions.
  • Importance of Motion Prior Initialization: Bootstrapping from the single-agent motion prior is essential for convergence. Without it, the RL process fails completely (0% success rate) or succumbs to reward hacking—exploiting the objective function to produce unnatural, non-functional motions. (Note: Entries marked with in Table 3 denote these reward-hacking failures).
@@
  • Enhanced Stability in Bed-based Assistance: We observe that our method yields more stable assistance, as measured by a COM stability metric (Appendix C).
Figure 5: Qualitative results on HHI-Assist. Red boxes indicate correct hand adjustment and support.
@@ 4.2 Generalist Policy Evaluation
Table 4 summarizes the tracking accuracy for generalist policies evaluated on the Inter-X dataset. When trained directly on a diverse subset of 30 interaction clips, AssistMimic achieves a success rate of 77.3%, making it, to our knowledge, the first physics-based method to handle such a broad range of tightly coupled interactions. Furthermore, we observe that distilling specialist policies into a single generalist agent via DAgger yields substantial performance gains. The top two rows of Figure LABEL:fig:teaser present qualitative results on the Inter-X dataset obtained with a single generalist policy trained with AssistMimic. For additional qualitative results, please see Appendix B.
@@ 4.4 Tracking of Generated Interactions
Figure 6 illustrates the results of applying AssistMimic to trajectories synthesized by a text-conditioned motion diffusion model [25] (details in Appendix F). The diffusion model was trained on ”Help-up” interactions from the Inter-X dataset. During inference, we prompt it with unseen text descriptions from the same category. While the generated kinematics capture the interaction semantics, they often contain physical artifacts such as foot sliding and penetration. AssistMimic successfully converts dense interaction kinematics into physically plausible motions. This result demonstrates that AssistMimic is effective not only for real interaction data but also for generated data, highlighting its generality and broad applicability.
@@
(a) Generalization to unseen interactions.
(b) Failure cases due to limited hand dexterity.
Figure 7: Qualitative results: unseen interactions and failures.
@@ 4.5 Generalization to Unseen Supportive Actions
As shown in Figure 7 (a), AssistMimic also generalizes to unseen interaction categories from Inter-X, such as support-by-arm. Even though this interaction pattern is not observed during training, the humanoid successfully follows the motion while placing a supportive hand on the partner’s arm and avoiding body collisions.
4.6 Failure cases and limitations.
Figure 7 (b) shows failures in motions that require precise grasping and lifting of the recipient’s arms. These failures are primarily caused by limited hand dexterity, as grasping a human arm requires fine, coordinated finger control that is difficult to learn from noisy demonstrations and are further exacerbated by the recipient’s weight. Tighter integration between motion planning and tracking is needed for more adaptive coordination. Incorporating visual observations is necessary to improve robustness to dynamic partner states. Transferring the learned behaviors from simulation to real humanoid systems remains an important challenge.
@@ 5 Conclusion
We introduced AssistMimic, a multi-agent reinforcement learning framework for learning physics-aware, tracking controllers that reproduce and adapt to human–human interactions, including assistive behaviors. Unlike prior approaches that treat interaction partners independently or rely on fixed replayed trajectories, AssistMimic jointly learns policies for both agents, enabling reactive and physically consistent coordination. AssistMimic improves exploration and stability in RL through three key innovations: (1) weight initialization from single-human tracking controllers, which stabilizes early learning; (2) dynamic reference retargeting, allowing the policy to adjust to the recipient’s current pose; and (3) contact-promoting rewards, encouraging physically meaningful interactions. Experiments on the Inter-X and HHI-Assist show that AssistMimic is the first method capable of successfully tracking assistive interaction sequences. More broadly, our results demonstrate a generalizable framework for extending single-agent imitation learning to physically grounded multi-agent domains, with promising implications for assistive robotics, collaborative manipulation, and embodied simulation of social interactions.
@@ Appendix B Additional Qualitative comparison
Figure 8 visualizes the ablation results of the specialist policies on the Inter-X dataset. Each row corresponds to a different variant, and frames are shown from left to right in chronological order. The first and second rows compare the best AssistMimic policy (w/o dynamic reference retargeting) and the variant without the contact-promoting reward on the same help-up motion. The third row shows two separate examples for the variant without weight initialization, because the humanoids quickly lose balance and the episodes terminate early.
@@
Table 5: Recipient COM standard deviation ($\downarrow$)
Model Seen Mass $\times 1.5$ Max hip $\tau$ $\times 0.5$
Ours 0.0921 0.0738 0.0865
(-) Dyn Retar 0.1038 0.0902 0.0924
(-) Cont 0.0938 0.0838 0.0849
Appendix C COM Stability Metric
To quantify the stability of assistive interactions, we measure the temporal variation of the recipient’s center of mass (COM) during each episode.
Definition.
Let $\mathbf{c}_{t}\in\mathbb{R}^{3}$ denote the 3D position of the recipient’s COM at timestep $t$, computed from the mass-weighted average of all body segments. We define the COM stability metric as the standard deviation of $\mathbf{c}_{t}$ over time:
$\sigma_{\text{COM}}=\sqrt{\frac{1}{T}\sum_{t=1}^{T}\left\|\mathbf{c}_{t}-\bar{\mathbf{c}}\right\|^{2}},\qquad\bar{\mathbf{c}}=\frac{1}{T}\sum_{t=1}^{T}\mathbf{c}_{t},$ (13)
where $T$ is the number of timesteps in the episode.
Evaluation protocol.
To ensure a fair comparison, we compute $\sigma_{\text{COM}}$ only on motion sequences where all compared methods successfully complete the task. This avoids bias due to early termination or failure cases, which would otherwise result in shorter trajectories and artificially lower variance.
Interpretation.
A lower $\sigma_{\text{COM}}$ indicates more stable assistance, as the recipient’s body is maintained with less oscillation or unintended movement. This metric is particularly relevant for bed-based caregiving scenarios (e.g., HHI-Assist), where maintaining a steady posture is critical.
Results.
As summarized in Table 5, AssistMimic achieves lower recipient COM standard deviation compared to the ablated variants on the HHI-Assist when evaluated on motions successfully completed by all policies. Together with the results in Table 3, this indicates that our approach achieves both high success rates and stable assistance.
Appendix D Additional implementation details
@@
For the multi-agent RL optimization, we use Proximal Policy Optimization (PPO) [21] for both the supporter and the recipient policies. For the specialist comparisons in Table 2, 3, policies are trained for 2k iterations, which we found sufficient for these focused subsets. For the generalist evaluation in Table 4, our AssistMimic w/o DAgger (1st line) is trained for $9\text{k}$ iterations. Similarly, we train specialist teacher policies for $9\text{k}$ iterations to provide high-fidelity supervision, from which the final generalist policy is distilled via DAgger [18] over $4\text{k}$ epochs. The learning rate for both agents is set to $5\times 10^{-6}$, and we apply a step decay by a factor of $0.1$ once training reaches $600$ PPO iterations. The critic network is implemented as an MLP that takes the same input as the actor, augmented with a role label indicating whether the agent is the recipient or the caregiver (supporter).
@@
In many HHI-Assist bed sequences, the supporter steps away from the recipient near the end of the motion after the assistance is completed. Because the recipient is no longer supported during this phase, falls occurring in this stage should not be considered assistive failures. Thus, for quantitative evaluations (Table 3, Table 5), we discard the final 15 frames of each 30 FPS motion clip and compute all metrics on the remaining frames.
@@
Appendix E Detailed reward breakdown